Show your work. Express the operands in IEEE single precision number format and

perform each operation below. Store the final result of each operation back in IEEE single

precision number format.

a.1290.25 + 87.125

b. 96.50 - 31.25

Please answer both
Answer #1

a.

1290.25 = 1.260009765625 * 2^(10)

87.125 = 1.361328125 * 2^(6)

1290.25 + 87.125 = 1377.375

1377.375 = 1.3450927734375 * 2^(10)

b.

96.50 = 1.5078125 * 2^(6)

31.25 = 1.953125 * 2^(4)

96.50 - 31.25 = 65.25

65.25 = 1.01953125 * 2^(6)